The length of a rectangle is four times its width.
If the area of the rectangle is 324ft^2 , find its perimeter.

Answer by scott8148(6628)

w * 4w = 324 ft^2

4w^2 = 324 ft^2 ___ w^2 = 81 ft^2 ___ w = 9 ft

l = 4w = 4(9) = 36 ft

perimeter = 36 + 36 + 9 + 9 = 90 ft